A wine company needs to blend a California wine with a 5% alcohol content and a french wine with a 9% alcohol content to obtain 200 gallons of wine with a 7% alcohol content. How many gallons of each kind of wine must be used?
Answer by mananth(16946)

----------------- percent ----------------quantity
Californian wine 5 ----------------x gallons
French wine 9 ----------------200-x gallons
Mixture 7.00% ----------------200

5x+9(200-x)=7*200

5x+1800 -9x=1400
5x-9x=1400-1800
-4x= -400
/ -4
x=100 gallons 5.00% Californian wine
100 gallons 9.00% French wine
m.ananth@hotmail.ca